OpenStudy (anonymous):

A number is said to be a palindrome if it reads the same backwards as forwards(e.g.58485). How many 5-digit numbers are palin-dromes?

OpenStudy (anonymous):

2*9C2 = 2*9*8=144 ( might be so)

OpenStudy (anonymous):

The numbers will have the following pattern: abcba where a, b, c are digits. since it's a 5 digit number: a = 1-9 (9 digits possible) b= 0-9 (10 digits possible) c= 0-9 (10 digits possible) number of possibilities are 9*10*10 = 900

OpenStudy (anonymous):

a cannot be 0 , coz then the number will have a leading 0 as a digit making it a 4 digit number.

OpenStudy (anonymous):

thanks.